# Doracamom: Joint 3D Detection and Occupancy Prediction with Multi-view 4D Radars and Cameras for Omnidirectional Perception

### v2

@@ IV-B Implementation Details
For the OmniHD-Scenes dataset, we constrain the point cloud range to $(-60,60)$ m, $(-40,40)$ m, and $(-3,5)$ m along the X-, Y-, and Z-axes, respectively. The radar input consists of 3-frame accumulated point clouds, with each point characterized by a feature vector $[x,y,z,Power,SNR,v_{xr},v_{yr}]$, where $Power$ and $SNR$ represent amplitude and signal-to-noise ratio, while $v_{xr}$ and $v_{yr}$ denote the compensated absolute velocity components. All six camera images are resized to $544\times 960$. The low-resolution voxel query dimensions $H_{V}\times W_{V}\times Z_{V}$ are set to $80\times 120\times 8$, the BEV feature map size $H\times W$ is $160\times 240$, and the occupancy ground truth resolution is $160\times 240\times 16$. We employ a DETR-based detection head with 900 object queries and retain the top 300 predicted boxes with highest confidence scores during inference. For both VoD and TJ4DRadSet datasets, we followed exact same settings as most of the state-of-the-arts [68, 70, 71].
@@
All ablation experiments are conducted on the OmniHD-Scenes dataset under a multi-task learning setting with two temporal frames, unless otherwise specified.
@@ IV-C Performance Comparison with State-of-the-arts
Figure 6: Qualitative results of Doracamom on the OmniHD-Scenes test set. The left column shows predicted 3D bounding boxes projected onto six camera views, where cars, pedestrians, riders, and large vehicles are displayed in yellow, blue, green, and pink, respectively. The middle column presents the BEV perspective, with black points representing LiDAR data and red points indicating 4D radar data, where ground truth and predicted boxes are shown in green and blue. The right column shows the predicted occupancy grids and ground truth occupancy grids, using the same color scheme as Table II. These results demonstrate that Doracamom achieves robust multi-task perception performance across diverse scenarios. Please zoom in to see more detail.
Figure 8: Qualitative results of Doracamom on the TJ4DRadSet and VoD datasets. For each sample, the upper row shows predicted 3D bounding boxes projected onto camera views, and the lower row presents BEV visualization with red points representing 4D radar data and blue boxes indicating detection results.
@@ IV-C1 Results on OmniHD-Scenes
TABLE I: Comparison of state-of-the-art approach with ours, for 3D object detection task on the OmniHD-Scenes test set. “L” denotes LiDAR, “C” denotes camera and “R” denotes 4D radar. ”-128” and ”-32” represent 128 lines and manually downsampled 32 lines, respectively. The last four columns show the AP for each object type. “Ped.” and “LVeh.” represent pedestrian and large vehicle, respectively. Bold and underline denote the first and the second best performances among all the approaches without using LiDAR. Doracamom-S does not utilize temporal information.
[…]
PointPillars-128 [3] (CVPR 2019) - L - 61.15 55.54 0.2825 0.1980 0.5223 1.8763 85.43 28.63 69.18 61.36
PointPillars-32 [3] (CVPR 2019) - L - 57.24 52.66 0.3040 0.2038 0.5692 1.8731 82.74 24.22 65.86 56.16
PointPillars [3] (CVPR 2019) - R - 23.82 37.21 0.6752 0.2447 0.3776 0.6789 52.74 0.69 28.57 13.29
RadarPillarNet [68] (IEEE T-IM 2023) - R - 24.88 37.81 0.6597 0.2389 0.3736 0.6982 52.99 2.06 29.45 15.02
LSS-Depth [20] (AAAI 2023) 544×960 C R50 22.44 26.01 1.0238 0.2230 0.5942 2.0138 50.42 4.37 24.56 10.42
BEVformer [21] (ECCV 2022) 544×960 C R50 26.49 28.10 1.1430 0.2315 0.5799 1.6666 50.74 11.69 30.42 13.11
BEVformer-T [21] (ECCV 2022) 544×960 C R50 29.17 30.54 1.1046 0.2346 0.4889 1.0797 53.64 14.48 33.55 15.01
BEVformer [21] (ECCV 2022) 864×1536 C R101-DCN 30.10 30.55 1.0633 0.2266 0.5331 1.6625 55.40 14.05 34.24 16.70
BEVformer-T [21] (ECCV 2022) 864×1536 C R101-DCN 32.22 32.57 1.0637 0.2271 0.4558 1.0683 57.61 17.37 38.02 15.87
PanoOcc [35] (CVPR 2024) 544×960 C R50 29.17 28.55 1.1500 0.2446 0.6378 1.6066 51.58 15.82 35.02 14.26
BEVFusion [80] (NeurIPS 2022) 544×960 C&R R50 33.95 43.00 0.5730 0.2165 0.3814 0.7474 56.25 11.66 50.90 16.99
RCFusion [68] (IEEE T-IM 2023) 544×960 C&R R50 34.88 41.53 0.5676 0.2135 0.3711 0.9208 57.17 12.87 51.35 18.11
Doracamom-S (ours) 544×960 C&R R50 37.60 41.31 0.6724 0.2329 0.4359 0.8579 58.94 17.84 52.72 20.89
Doracamom (ours) 544×960 C&R R50 39.12 46.22 0.6646 0.2331 0.3545 0.6151 61.12 19.83 53.35 22.18
@@
3D Object Detection. Table I presents the performance comparison of different methods on the OmniHD-Scenes test set for 3D detection tasks. Our proposed Doracamom achieves superior overall performance (39.12 mAP & 46.22 ODS) compared to other approaches based on 4D radar, camera, or their fusion. Specifically, it outperforms BEVFusion [80] by +5.17 mAP and +3.22 ODS, while surpassing RCFusion [68] by +4.24 mAP and +4.69 ODS. Even in the single-frame setting without the DTE module, our model outperforms all other methods in terms of mAP. Furthermore, Doracamom significantly narrows the performance gap with LiDAR-based PointPillars [3] (46.22 ODS vs. 55.54 ODS), which demonstrates both the effectiveness of our proposed architecture and the tremendous potential of low-cost sensor configurations in autonomous driving perception systems. In terms of TP metrics, our method achieves best performance in both mAOE and mAVE, reaching 0.3545 and 0.6151, respectively. Notably, benefiting from the Doppler information inherent in 4D radar, approaches leveraging 4D radar independently or incorporating 4D radar-camera fusion demonstrate remarkable velocity estimation accuracy compared to camera-only or LiDAR-based methods, exhibiting substantially lower errors in velocity predictions. Furthermore, Doracamom achieves even more precise velocity estimation through efficient exploitation of 4D radar features and the DTE module. For each object category, our method achieves the highest AP scores across all classes: Car, Pedestrian, Rider, and Large Vehicle. Notably, all methods exhibit relatively low detection accuracy for pedestrians and large vehicles, which can be attributed to several challenging factors in the dataset. The dataset contains numerous crowded scenes with dozens of people in small areas, leading to severe occlusions. Moreover, the features extracted from both image and radar data are often incomplete and less distinctive. Additionally, pedestrians occupy only a few grids in the BEV features, further increasing the detection difficulty. For Large Vehicles, their substantial size often results in incomplete contours in both radar and camera views, leading to significant size discrepancies.
The visualization results shown in Fig. 6 denote that Doracamom can provide reliable performance across both day and night scenarios. It achieves high detection accuracy in crowded and complex scenes, with only occasional missed detections of distant occluded objects. Fig. 7 illustrates the BEV feature maps of different methods. It can be observed that the feature map of Doracamom displays distinct object boundaries and highly distinguishable features, with no significant issues such as severe stretching or distortion of the objects.
@@
BEVFormer [21] (ECCV 2022) 544×960 C R50 27.04 14.97 20.64 5.87 14.40 16.68 1.52 3.64 20.64 46.61 16.19 14.80 3.69
BEVFormer-T [21] (ECCV 2022) 544×960 C R50 28.42 16.23 22.73 5.45 14.70 18.21 3.09 3.87 21.54 48.15 17.58 17.77 5.48
@@
BEVFormer [21] (ECCV 2022) 864×1536 C R101-DCN 28.30 16.41 23.72 6.37 16.33 20.44 1.78 3.78 22.21 48.55 17.88 15.49 3.99
BEVFormer-T [21] (ECCV 2022) 864×1536 C R101-DCN 29.74 17.49 24.90 6.48 16.45 21.49 2.87 4.62 22.51 49.92 18.59 18.53 5.96
@@
3D Semantic Occupancy. Table II presents the performance comparison of different methods on the OmniHD-Scenes validation set for occupancy prediction tasks. Our proposed Doracamom achieves superior overall performance (33.96 SC IoU & 21.81 mIoU) compared to other approaches. When BEVFormer [21] utilizes a larger backbone network (R101-DCN) and higher resolution image input (864*1536), its performance surpasses multi-sensor fusion methods like M-CONet [81] that combines camera and 4D radar data. Nevertheless, with our well-designed architecture, even without using temporal information, Doracamom-S significantly outperforms BEVFormer-T by +1.72 SC IoU and +2.00 mIoU. Moreover, examining the IoU metrics for each category, we observe that both Doracamom and Doracamom-S achieve superior performance compared to other models in detecting foreground objects of particular interest for object detection, specifically cars, pedestrians, riders, and large vehicles. This performance advantage demonstrates that in a multi-task setting, object detection and occupancy prediction can work synergistically to facilitate both tasks. This significant advantage is also visually apparent in Fig. 6, showing that Doracamom maintains strong occupancy prediction performance in both day and night scenarios. In some cases, it even successfully completes ground holes present in the ground truth. However, the model still exhibits some prediction errors in occluded regions.
@@
TABLE IV: Comparison of resources consumption and efficiency with other state-of-the-arts (tested on the 4090D GPU).
@@
Doracamom-2frames 4.72G 52.67M 4.4 38.21 44.52 33.16 21.44
Resources Consumption and Efficiency. Table IV presents a comparison of different models in terms of resource consumption and efficiency. Compared to existing methods, our Doracamom series models demonstrate an excellent balance between performance and efficiency. In terms of resource consumption, Doracamom-S requires only 4.71G memory and 49.63M parameters, making it more lightweight than BEVFusion ($\sim$8G memory and 57M parameters) and PanoOcc (5.03G memory and 51.94M parameters). Even with 2 frames incorporated, Doracamom-2frames maintains relatively low resource usage (4.72G memory, 52.67M parameters). Regarding inference efficiency, Doracamom-S and Doracamom-2frames achieve 4.8FPS and 4.4FPS respectively, significantly outperforming the BEVFusion series (3.2-3.6FPS). While slightly slower than PanoOcc (5.5FPS), our models demonstrate substantial performance advantages: Doracamom-2frames achieves optimal performance across all evaluation metrics, significantly surpassing other methods.
@@ IV-C2 Results on VoD & TJ4DRadSet
TABLE V: Comparison of state-of-the-art approach with ours, for 3D object detection task on VoD validation set and TJ4DRadSet test set, respectively. “C” denotes camera and “R” denotes 4D radar. indicates the use of extra LiDAR data during training procedure. Bold and underline denote the first and the second best performances among all the approaches. To ensure fair comparison, Doracamom-S does not utilize temporal information, same as others.
@@
Fig. 8 presents visualization results from both TJ4DRadSet and VoD datasets. The results show that Doracamom-S maintains excellent performance across diverse scenarios, demonstrating low localization error compared to ground truth and accurate category prediction. Only occasional missed detections are observed under challenging occlusion conditions.
IV-D Ablation Study
@@
Ours w/o Aux. 37.84 44.42 32.81 21.30
Ours w/ Aux. 38.21 44.52 33.16 21.44
Ablation Study on CMF. Table IX presents the ablation study results of the Cross-modal BEV-Voxel Fusion (CMF) module. We compare CMF with two basic feature fusion strategies (feature addition and concatenation). Our proposed CMF module, which adaptively fuses heterogeneous features from BEV and voxel spaces through attention-weighted mechanisms, demonstrates clear advantages even without auxiliary tasks (Ours w/o Aux.): achieving 0.19 improvement in mIoU compared to feature addition, and 0.25 improvement in ODS compared to feature concatenation. With the introduction of auxiliary tasks (Ours w/ Aux.) - foreground/background BEV segmentation and occupied/non-occupied binary prediction - the model achieves comprehensive performance improvements (+0.37 mAP, +0.10 ODS, +0.35 IoU, +0.14 mIoU) over the version without auxiliary tasks. These experimental results not only validate the effectiveness of our adaptive fusion strategy in integrating multi-modal features but also demonstrate that auxiliary tasks can successfully enhance feature representation through additional supervisory signals.
V Conclusion

### v3

@@ IV-B Implementation Details
For the OmniHD-Scenes dataset, we constrain the point cloud range to (60,60)(-60,60) m, (40,40)(-40,40) m, and (3,5)(-3,5) m along the X-, Y-, and Z-axes, respectively. The radar input consists of 3-frame accumulated point clouds, with each point characterized by a feature vector $x,y,z,Power,SNR,v_{xr},v_{yr}]$, where $Power$ and $SNR$ represent amplitude and signal-to-noise ratio, while $v_{xr}$ and $v_{yr}$ denote the compensated absolute velocity components. All six camera images are resized to $544\times 960$. The low-resolution voxel query dimensions $H_{V}\times W_{V}\times Z_{V}$ are set to $80\times 120\times 8$, the BEV feature map size $H\times W$ is $160\times 240$, and the occupancy ground truth resolution is $160\times 240\times 16$. We employ a DETR-based detection head with 900 object queries and retain the top 300 predicted boxes with highest confidence scores during inference. For both VoD and TJ4DRadSet datasets, we followed exact same settings as most of the state-of-the-arts [84, 67, 1].
@@
All ablation experiments are conducted on the OmniHD-Scenes dataset under a multi-task learning setting with 2 temporal frames, unless otherwise specified.
@@ IV-C Performance Comparison with State-of-the-arts
TABLE I: Comparison of state-of-the-art approach with ours, for 3D object detection task on the OmniHD-Scenes test set. “C” denotes camera and “R” denotes 4D radar. The last four columns show the AP for each object type. “Ped.” and “LVeh.” represent pedestrian and large vehicle, respectively. Bold and underline denote the first and the second best performances among all the approaches. Doracamom-S does not utilize temporal information. All FPS metrics are tested on a single NVIDIA GeForce RTX 4090D. † denotes the end-to-end inference speed for jointly performing both 3D detection and occupancy prediction tasks.
[…]
PointPillars [29] (CVPR 2019) - R - 23.82 37.21 0.6752 0.2447 0.3776 0.6789 52.74 0.69 28.57 13.29 62.2
RadarPillarNet [84] (IEEE T-IM 2023) - R - 24.88 37.81 0.6597 0.2389 0.3736 0.6982 52.99 2.06 29.45 15.02 60.3
LSS-Depth [32] (AAAI 2023) 544×960 C R50 22.44 26.01 1.0238 0.2230 0.5942 2.0138 50.42 4.37 24.56 10.42 3.7
BEVDet4d [22] (arXiv 2022) 544×960 C R50 25.22 28.66 0.9344 0.2352 0.5470 1.1881 52.07 8.07 31.92 8.86 18.3
BEVFormer-S [33] (ECCV 2022) 544×960 C R50 26.49 28.10 1.1430 0.2315 0.5799 1.6666 50.74 11.69 30.42 13.11 11.4
BEVFormer [33] (ECCV 2022) 544×960 C R50 29.17 30.54 1.1046 0.2346 0.4889 1.0797 53.64 14.48 33.55 15.01 11.4
SparseBEV [38] (ICCV 2023) 544×960 C R50 30.14 30.38 1.0272 0.2216 0.5534 1.2188 54.68 17.28 34.52 14.08 4.9
BEVFormer-S [33] (ECCV 2022) 864×1536 C R101-DCN 30.10 30.55 1.0633 0.2266 0.5331 1.6625 55.40 14.05 34.24 16.70 4.7
BEVFormer [33] (ECCV 2022) 864×1536 C R101-DCN 32.22 32.57 1.0637 0.2271 0.4558 1.0683 57.61 17.37 38.02 15.87 4.7
PanoOcc [62] (CVPR 2024) 544×960 C R50 29.17 28.55 1.1500 0.2446 0.6378 1.6066 51.58 15.82 35.02 14.26 5.5
BEVFusion [34] (NeurIPS 2022) 544×960 C&R R50 33.95 43.00 0.5730 0.2165 0.3814 0.7474 56.25 11.66 50.90 16.99 3.6
RCFusion [84] (IEEE T-IM 2023) 544×960 C&R R50 34.88 41.53 0.5676 0.2135 0.3711 0.9208 57.17 12.87 51.35 18.11 3.6
RCBEVDet [37] (CVPR 2024) 544×960 C&R R50 35.53 45.41 0.5138 0.2305 0.3614 0.6825 62.35 10.11 54.60 15.06 5.2
Doracamom-S (ours) 544×960 C&R R50 37.60 41.31 0.6724 0.2329 0.4359 0.8579 58.94 17.84 52.72 20.89 4.8
Doracamom (ours) 544×960 C&R R50 39.12 46.22 0.6646 0.2331 0.3545 0.6151 61.12 19.83 53.35 22.18 4.2
Figure 6: Qualitative comparison for 3D object detection on the OmniHD-Scenes test set. The left column shows our 3D bounding box predictions projected onto the six camera views. The right column shows the comparison between RCFusion [84], RCBEVDet [37], our method, and the ground truth (GT). In the BEV diagrams, LiDAR points are represented in black, and 4D radar points in red. The highlighted regions indicate areas where our method demonstrates clear advantages in reducing false positives and missed detections compared to the other methods.
@@ IV-C1 Results on OmniHD-Scenes
3D Object Detection. Table I presents the performance comparison of different methods on the OmniHD-Scenes test set for 3D object detection. Our proposed Doracamom achieves superior overall performance (39.12 mAP & 46.22 ODS) compared to other approaches based on 4D radar, camera, or their fusion. Specifically, it surpasses the recent state-of-the-art method RCBEVDet [37] by +3.59 mAP and +0.81 ODS. The performance gap is even larger when compared to other strong baselines like BEVFusion [34] (+5.17 mAP, +3.22 ODS) and RCFusion [84] (+4.24 mAP, +4.69 ODS). Even in the single-frame setting, our Doracamom-S outperforms all other methods in terms of mAP. In terms of TP metrics, our method achieves best performance in both mAOE and mAVE, reaching 0.3545 and 0.6151, respectively. Notably, benefiting from the Doppler information inherent in 4D radar, approaches leveraging 4D radar independently or incorporating 4D radar-camera fusion demonstrate remarkable velocity estimation accuracy compared to camera-only methods, exhibiting substantially lower errors in velocity predictions. Furthermore, Doracamom achieves even more precise velocity estimation through efficient exploitation of 4D radar features and the DTE module. For each object category, our method achieves the highest or second-highest AP scores across all classes. Notably, all methods exhibit relatively low detection accuracy for pedestrians and large vehicles, which can be attributed to several challenging factors in the dataset. The dataset contains numerous crowded scenes with dozens of people in small areas, leading to severe occlusions. Moreover, the features extracted from both image and radar data are often incomplete and less distinctive. Additionally, pedestrians occupy only a few grids in the BEV features, further increasing the detection difficulty. For Large Vehicles, their substantial size often results in incomplete contours in both radar and camera views, leading to significant size discrepancies.
Fig. 6 provides a qualitative comparison of Doracamom with other methods on the 3D detection task. The figure shows the performance of our model in various weather conditions, including daytime, nighttime, and rain in complex urban traffic scenes. It is clear that our Doracamom shows superior performance compared to RCFusion and RCBEVDet. The highlighted regions show that our method reduces both missed detections and false positives when dealing with dense objects and distant objects, proving its stronger robustness and generalization ability. Fig. 8 illustrates the BEV feature maps of different methods. It can be observed that the feature map of Doracamom displays distinct object boundaries and highly distinguishable features, with no significant issues such as severe stretching or distortion of the objects.
@@
BEVFormer-S [33] (ECCV 2022) 544×960 C R50 27.04 14.97 20.64 5.87 14.40 16.68 1.52 3.64 20.64 46.61 16.19 14.80 3.69
BEVFormer [33] (ECCV 2022) 544×960 C R50 28.42 16.23 22.73 5.45 14.70 18.21 3.09 3.87 21.54 48.15 17.58 17.77 5.48
@@
BEVFormer-S [33] (ECCV 2022) 864×1536 C R101-DCN 28.30 16.41 23.72 6.37 16.33 20.44 1.78 3.78 22.21 48.55 17.88 15.49 3.99
BEVFormer [33] (ECCV 2022) 864×1536 C R101-DCN 29.74 17.49 24.90 6.48 16.45 21.49 2.87 4.62 22.51 49.92 18.59 18.53 5.96
@@
Figure 7: Qualitative comparison for semantic occupancy prediction on the OmniHD-Scenes test set. The left column shows the multi-view camera images for reference. The right columns compare the predicted occupancy of SurroundOcc [63], M-CONet [60], our method, and the ground truth (GT), using the same color scheme as Table II. The highlighted regions emphasize areas where our method demonstrates superior performance in capturing fine-grained details and achieving higher completeness, especially in challenging scenarios.
@@
3D Semantic Occupancy. Table II presents the performance comparison of different methods on the OmniHD-Scenes test set for occupancy prediction tasks. Our proposed Doracamom achieves superior overall performance (33.96 SC IoU & 21.81 mIoU) compared to other approaches. When BEVFormer [33] utilizes a larger backbone network (R101-DCN) and higher resolution image input (864×1536), its performance surpasses multi-sensor fusion methods like M-CONet [60] that combines camera and 4D radar data. Nevertheless, with our well-designed architecture, even the Doracamom-S which does not utilize temporal information, that is, without using historical BEV and voxel features, significantly outperforms BEVFormer by +1.72 SC IoU and +2.00 mIoU. Moreover, examining the IoU metrics for each category, we observe that both Doracamom and Doracamom-S achieve superior performance compared to other models in detecting foreground objects of particular interest for object detection, specifically cars, pedestrians, riders, and large vehicles. This performance advantage demonstrates that in a multi-task setting, object detection and occupancy prediction can work synergistically to facilitate both tasks. This advantage is also visually demonstrated in the qualitative comparison in Fig. 7. The figure compares our method with SurroundOcc, M-CONet, and the ground truth (GT). The results show that Doracamom is superior to the other methods in terms of prediction completeness and capturing fine-grained details in both daytime and nighttime scenarios.
@@
TABLE IV: Comparison of resources consumption and efficiency with other state-of-the-art fusion methods. (tested on the 4090D GPU).
@@
Doracamom 6.04G 53.30M 4.2 39.12 46.22 33.96 21.81
Resources Consumption and Efficiency.
Table IV provides a comprehensive comparison of various fusion methods in terms of the resource consumption and inference efficiency. It’s crucial to note that the FPS of our Doracamom method represents the speed while performing both 3D object detection and semantic occupancy prediction at the same time. Compared to other methods, the Doracamom series exhibits excellent lightweight performance. Doracamom-S has the lowest parameters (49.63M), and its GPU memory usage is only slightly more than RCBEVDet. In the trade-off between efficiency and performance, Doracamom achieves 4.2 FPS, which is close to the 5.2 FPS of RCBEVDet’s single-task operation. However, Doracamom achieves this while handling two tasks simultaneously and delivering SOTA performance of 46.22 ODS and 21.81 mIoU. Compared to BEVFusion, our method has a clear advantage in both speed and accuracy. This highlights that our unified framework finds a better balance between computational efficiency and performance.
As for deployment, the current Doracamom speed is measured under an unoptimized PyTorch framework. However, we’re confident that we could significantly improve the model’s speed with standard industrial optimizations, like hardware acceleration, quantization, pruning, and resolution compression. Our future work will focus on exploring these optimizations to meet deployment requirements.
@@ IV-C2 Results on VoD & TJ4DRadSet
TABLE V: Comparison of state-of-the-art approach with ours, for 3D object detection task on VoD validation set and TJ4DRadSet test set, respectively. “C” denotes camera and “R” denotes 4D radar. indicates the use of extra LiDAR data during training procedure. Bold and underline denote the first and the second best performances among all the approaches. To ensure fair comparison, Doracamom-S does not utilize temporal information.
@@ IV-D Ablation Study
CMF 38.21 44.52 33.16 21.44
TABLE X: Ablation study on extrinsic parameter perturbation.
Max Perturbation Metrics
[…]
0.0 0.0 38.21 44.52 33.16 21.44
$\pm 2.0$ $\pm 0.2$ 38.14 44.48 32.98 21.34
$\pm 5.0$ $\pm 0.5$ 37.80 44.23 32.23 20.87
$\pm 10.0$ $\pm 1.0$ 36.56 43.42 30.22 19.62
Ablation Study on CMF. Table IX presents the ablation study results of the CMF module. We compared CMF with two basic feature fusion strategies, feature addition and concatenation. The results demonstrate that our CMF module outperforms all other methods across all metrics. In comparison to the simple addition strategy, our CMF enhances mAP by +0.49 and mIoU by +0.33. Against the concatenation strategy, the benefits of our method are even more evident, with improvements of +0.64 in mAP and +0.60 in mIoU. These results strongly validate the effectiveness of our CMF module design. By combining adaptive attention mechanisms with auxiliary supervision signals, it can more effectively integrate heterogeneous features from the BEV and voxel spaces than simpler fusion methods, thus generating more discriminative and comprehensive feature representations for downstream tasks.
TABLE XI: Ablation study on loss components.
Loss Components Metrics
[…]
37.16 43.21 - -
37.84 44.42 32.81 21.30
38.05 44.67 32.80 21.35
38.21 44.52 33.16 21.44
TABLE XII: Ablation study on encoder layers.
Layers Metrics
[…]
1 35.84 43.52 32.57 20.60
2 37.06 44.34 32.54 20.77
3 38.21 44.52 33.16 21.44
4 38.51 44.55 32.59 21.27
Ablation study on extrinsic parameter perturbation. Table X shows the robustness evaluation results of Doracamom when the extrinsic parameters of cameras are disturbed. To simulate real-world calibration drift, we manually added random noise to the translation and rotation parameters of cameras. The noise was sampled independently from a symmetric uniform distribution of $\delta,+\delta]$, where $\delta$ is the maximum error. The results clearly show that he performance of Doracamom decreases smoothly as the noise level increases, without any sudden collapse. When a small noise is introduced ($\pm 2.0cm$, $\pm 0.2^{\circ}$), the performance of our model is almost unaffected, with only 0.04 and 0.1 decrease in ODS and mIoU respectively. Even with larger noise ($\pm 10.0cm$, $\pm 1.0^{\circ}$), the model still maintained a relatively high performance level, with 43.42 ODS and 19.62 mIoU. This demonstrates the robustness of Doracamom against a certain range of calibration errors, which is crucial for practical deployment. In the future, we can make the model more robust by training with calibration noise or by adding a lightweight online self-calibration module.
Ablation study on loss components. Table XI shows the impact of different loss functions on the model performance. We first use a single-task model that only uses $\mathcal{L}_{det}$ for 3D object detection as the baseline, with a performance of 37.16 mAP and 43.21 ODS. When we add the $\mathcal{L}_{occ}$ to this baseline for multi-task training, mAP and ODS increase by 0.68 and 1.21, respectively. This demonstrates that adding the occupancy prediction task helps Doracamom learn richer spatial features, which positively affects the detection task. When we next add the BEV segmentation auxiliary loss $\mathcal{L}_{aux}^{seg}$, nearly all metrics improve. This indicates that guiding the model to better distinguish between foreground and background in the BEV space is beneficial for both 3D semantic understanding and object detection. when we further add the $\mathcal{L}_{aux}^{occupied}$, Doracamom achieves its highest scores for mAP, IoU, and mIoU. This shows that combining the two auxiliary losses guides the model to learn more comprehensive features, resulting in the best overall performance across both tasks.
Ablation study of encoder layers. Table XII shows ablation study results on the number of Transformer layers $L$ in the Voxel Queries Encoder. We observe that as $L$ increased from 1 to 3, most of the metrics show significant and consistent improvement. For instance, mAP increases from 35.84 to 38.21 and mIoU reaches at 21.44. This proves that a deeper encoder can effectively learn more fine-grained features. However, when $L$ increases to 4, the performance gain becomes very limited, and the metrics for occupancy prediction actually decrease. This indicates that the benefits of adding more depth are diminishing. Therefore, considering the trade-off between performance and computational efficiency, we choose $L=3$ as the final configuration for our model.
V Conclusion and Future Work